OpenStudy (stormswan):

please help, I will give a medal :) Solve for x: the quantity of x plus 4 over 3 = 2.

OpenStudy (shamil98):

\[\huge \frac{ x+4 }{ 3 } = 2\] You need to get rid of the 3, so multiply by 3 to both sides what do you get?

OpenStudy (stormswan):

wait to both sides like to the 2? then it would be 6...

OpenStudy (shamil98):

\[\huge \cancel 3 * \frac{ x+4 }{ \cancel3 } = 2*3\] \[\huge x+4 = 6\]

OpenStudy (shamil98):

I trust you can do it from here?

OpenStudy (stormswan):

yeah thats what I meant lol, and yeah I can, Thanks :)